### v1

# How Far Are We from Believable AI Agents? A Framework for Evaluating the Believability of Human Behavior Simulation

Abstract
Human behavior simulation of AI agents necessitates the agents to possess a quality of believability, which is crucial as it facilitates users in establishing trust toward the agents and streamlines the fulfillment of the agents’ goal. While recent advancements in Large Language Model (LLM) based agents have improved human behavior simulation, challenges inherent to LLMs (e.g., long context modeling) can undermine their believability. Consequently, evaluating AI agent believability becomes imperative. Unfortunately, prior research often neglects the negative impacts of LLM deficiencies. To address these gaps, we introduce two metrics for assessing LLM-based agent believability: consistency, and robustness, together with a benchmark, SimulateBench, with which, we evaluate the consistency and robustness of agents implemented with popular LLMs. We find that agents (i) struggle to accurately depict character information when presented with lengthy profile inputs; (ii) exhibit vulnerability to profile perturbations, and (iii) are significantly affected by certain key factors that impact their overall believability.  [^1] Code and SimulateBench are public at GitHub
1 Introduction
AI agents (Wooldridge and Jennings, 1995; Macal and North, 2005) have shown promise to simulate human behavior and social interaction, which can empower applications ranging across prototyping social theories (Aher et al., 2023b; Horton, 2023a; Kosinski, 2023; Kovac et al., 2023), generating synthetic research data (Hämäläinen et al., 2023; Wang et al., 2023b) and building non-player characters (Laird and VanLent, 2001). These applications necessitate the agents’ simulated human behavior to possess a quality of believability : the generated behavior enables others to suspend their disbelief towards the agents (Ortony et al., 2003), which is crucial because it facilitates others in establishing trust toward the agents and streamlines the fulfillment of the agents’ goals in these applications.
Figure 1: An illustrative example of the “Consistency”, and “Robustness”. Consistency measures whether the LLMs’ generated human behavior accurately depicts the identity information; Robustness measures whether the generated human behavior will be influenced by the perturbation in the profile.
Recently, the LLM-based agents (Park et al., 2023) have accelerated the realization of human behavior simulation by providing the LLMs with some identity information through prompting engineering (Liu et al., 2023b). Despite the initial results achieved now, some challenging issues of LLMs may pose a significant threat to the believability of LLM-based agents: (1) current LLMs do not perfectly make use of information in long input contexts (Liu et al., 2023a). This may impede LLM-based agents from capturing crucial information about characters in the long profile, which is usually integrated into the input to prompt agents to simulate humans. The profile may consist of personal data such as identity information, social roles (Wasserman and Faust, 1994), and relationships according to the simulation target. (2) LLMs lack robustness confronted with perturbed inputs (Perez and Ribeiro, 2022). Hence, minimal alterations to the profile input may result in significant fluctuations in the model’s performance within the same scenario, which will undermine the model’s believability. Unfortunately, the profile information usually undergoes dynamic alterations, such as a change in age, because agents are constantly evolving in the environment, which will cause the profile to update. Nevertheless, current studies (Park et al., 2023; Wang et al., 2023c; Shao et al., 2023; Cheng et al., 2023) fail to evaluate the believability along with the two issues. Firstly, the metrics proposed in these works for evaluation are not indicative of the agent’s believability when processing long and diverse profile input because these works only prompt limited identity information to the agents, which is of a concise nature. Secondly, there is little or no mechanism to thoroughly examine the potential impact of perturbations on the believability of LLMs and how these perturbations specifically affect their credibility.
To assess the negative impact brought by LLMs’ deficiency in processing long profile input and lack of robustness, we propose two metrics to evaluate the believability of LLM-based agents, as shown in Figure 1: (1) consistency: to what extent does the LLM-generated human behavior accurately depict the identity information, social roles, and relationships presented in the long profile input? (2) robustness: to what extent do the agents’ behaviors maintain robustness when faced with perturbations in the profile? To measure the consistency and robustness, we introduce SimulateBench: a framework for character data collection and evaluation of agents’ consistency and robustness ability. SimulateBench consists of three parts: the profile descriptive framework, the consistency dataset, and the robustness dataset. The profile descriptive framework is proposed to document the information about a person comprehensively. Based on the framework, we collect character datasets on characters for the agents to simulate. To measure the consistency, we assess whether the agent can correctly answer single-choice questions about the character that the agent is assigned to simulate in the consistency dataset. To measure the robustness, we perturb the profile of the assigned character and compare how the agent’s consistency ability changed faced with the profile perturbation.
Through the SimulateBench, we evaluate the level of believability of agents composed of API-based or open-source LLMs. Our findings show that: (1) agents perform poorly for consistency: they can not accurately depict the information in the long profile input, even if they are equipped with long context size; (2) agents exhibit a lack of robustness when faced with even nuanced profile perturbation. In further studies, we find that demographic factors, information position in the profile, and reasoning prompting will greatly influence the agents’ believability.
In summary, we propose two novel metrics of consistency and robustness to measure agents’ believability. To facilitate the assessment, we introduce the SimulateBench. We hope our work will inspire further research into the believability of agents’ human behavior simulation.
2 Related Work
2.1 LLM based Agent for Simulation
Recently, LLMs demonstrate intelligence comparable to humans (Rao et al., 2023a; Li et al., 2023; Jiang et al., 2023a; bench authors, 2023; Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2022). Many studies endeavor to harness the LLMs as a core brain in agents to simulate human behavior and social interactions (Maes, 1995; Wooldridge and Jennings, 1995; Xi et al., 2023). Park et al. (2022); Gao et al. (2023a) employs agents to simulate genuine human emotions, attitudes, and interaction behaviors in social networks. Park et al. (2023) applies in-context learning to enable the agent to emulate human behaviors in a sandbox world, such as organizing a party. Argyle et al. (2022) shows that the “algorithmic bias” within GPT-3 can be used to emulate response distributions from human subgroups. Horton (2023b) proposes LLMs can be used in behavioral economics experiments to pilot studies. Hämäläinen et al. (2023); Wang et al. (2023b) lets LLM-based agents simulate human behavior in a recommender system to gather user behavior data. In summary, LLM-based agents are being employed to simulate human behaviors and social interactions across several applications.
2.2 Evaluation for simulation of LLM-based Agents
Simulation of human behavior requires the agent to faithfully embody assigned roles and identities and proactively interact with other agents or humans (Franklin and Graesser, 1997; Elliott and Brzezinski, 1998; Laird and VanLent, 2001; Ortony et al., 2003; Wooldridge and Jennings, 1995; Jennings et al., 1998). Choi et al. (2023); Lin and Chen (2023); Fang et al. (2023); See et al. (2019) propose evaluations toward LLMs’ natural language understanding and generation. Rao et al. (2023b); Jiang et al. (2023b); Huang et al. (2023) evaluate LLMs’ ability to understand and maintain personality traits. Aher et al. (2023a) introduces the Turing Experiment to evaluate whether LLMs can simulate the behavior of a representative sample of participants in human subject research. Park et al. (2023); Gao et al. (2023a) propose a sandbox and an online social network to evaluate agents’ social interactions, respectively. However, to the best of our knowledge, no systematic and fine-grained benchmark exists to assess the agents’ believability. Hence, we aim to bridge this gap by proposing two novel metrics and constructing the first benchmark for simulation.
3 SimulateBench
To facilitate the measure of agents’ consistency and robustness, we introduce SimulateBench for characters’ profile collection and believability evaluation. We have collected the profiles of 56 characters and variants. There are 8400 questions in the consistency and robustness dataset. The statistic is shown in Table  1.
3.1 Profile Descriptive Framework and Datasets
Comprehensive profile information is necessary for agents to simulate human behavior accurately. Accordingly, we propose the profile descriptive framework and collect two character datasets based on this framework. For more details, please refer to the Appendix A.
Profile Descriptive Framework
Researchers in different fields pay attention to distinct profile information, given their different objectives. We propose a descriptive framework that documents a character’s information from three aspects: Immutable Characteristic, Social Role, Relationship. Immutable characteristic refers to characteristics that cannot be easily changed, such as name, gender, and age. Social role (Wasserman and Faust, 1994; Eagly and Wood, 2012) is conceptualized as a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation.
Character Dataset
We selected characters from TV dramas of popular genres [^2] https://www.imdb.com/list/ls023983860/: The Simpsons (Animated), Friends (Comedy), Breaking Bad (Crime), and The Rings of Power(Science fiction). According to the profile descriptive framework, we extract the profile information from the fandom [^3] https://www.fandom.com/. Human annotators carefully verify the collected data to ensure that the profiles are error-free. We also collected real people’s profile data to enhance the diversity of the dataset. To ensure privacy protection, we blur sensitive information, including name, age, and location.
Character Variants Dataset
To facilitate the measurement of LLMs’ robustness, we perturb the profile of characters in the character dataset by replacing the content of demographic factors: Education, Surname, Race, and Age. To prevent irrationality caused by the perturbation, a thorough examination of the consequences resulting from any modifications made to the initial profile is conducted. For instance, if Homer’s educational background is altered, any mention of it will likewise be adjusted.
3.2 Consistency Dataset
The consistency dataset is composed of single-choice questions. The agent needs to answer the questions in the dataset. According to the profile descriptive framework, there are three kinds of questions related to Immutable Characteristics, Social Roles, and Relationships. We manually generate the gold answers for every question and double-check the correctness of the answer. For every kind of question, we further categorize the questions into two classes according to their gold answer: Known and Not-Known. The gold answer of Not-Known is “There’s not enough information to answer this question.”. The Not-Known is the case that there is not enough information about the character in the profile for the agent to infer the answer.
3.3 Robustness Dataset
The robustness dataset is collected based on the consistency dataset. We have perturbed the character’s profile for each character in the character dataset to construct the character variants dataset. According to this perturbation, we modify the corresponding questions in the consistency dataset. Then, we include the modified questions in our robustness dataset. For instance, if we modify the age of Homer from 20 to 30, our initial step will involve duplicating the questions pertaining to Homer in the consistency dataset. Subsequently, we shall alter these questions and their gold answer to align with the age adjustment. After the alteration of these questions, we get the questions for Homer at the age of 30. All these altered questions for character variants make up the robustness dataset.
Statistical categories number
Characters 56
Avg tokens per profile 3277
Avg tokens per question 58
Avg questions per character #
Immutable Characteristic 41
Social Role 52
Relationship 57
Total benchmark questions 8400
Table 1: The statistics of SimulateBench. The tokens are counted with the embedding of GPT-4.
4 Simulation and Measuring
The two key aspects of believability are consistency and robustness. To measure the consistency and robustness ability of agents, we have three steps: (1) employing the agents to simulate human behavior of assigned character by prompt engineering; (2) measuring the consistency by requiring the agents to answer the single-choice questions in the consistency dataset; and (3) measuring robustness by perturbing the profile of the assigned character and comparing how the agent’s consistency ability changed faced with the profile perturbation.
4.1 Human Behavior Simulation
To construct an LLM-based Agent, the LLM is usually used as the core brain to process reasoning, memory recall, and other cognitive activities. So, LLM is crucial for the agents’ human behavior simulation. Given that, We employ LLM as an agent (Liu et al., 2023c) without integrating other modules to focus on the LLMs’ function for simulation444we use LLMs and LLM-based agents interchangeably.. To prompt the LLM to simulate human behavior, some factors are crucial: the instruction to explain how to simulate human behavior (I), the profile of specific characters (P), and some description of the task that the LLM needs to finish (T). Formally, the LLMs ($P$) are supposed to generate a response $\text{Y}=y_{1},\ldots,y_{n}$:
[…]
Blew will introduce the setting of I, P, and T.
I: Simulate Human Behavior
For models like GPT-3.5 that have gone through RLHF, (Wirth et al., 2017; Stiennon et al., 2020), the RLHF will equip LLMs with specific language preferences and habits, such as "as a language model", which will harm the believability of the agent. To overcome these issues, we set an instruction prompt template to instruct the LLM on how to simulate human behavior. Please refer to Appendix B for details of the prompt template.
P: Profile of Specific Characters
According to the Profile Descriptive Framework, we will fill in corresponding information in the instruction prompt template. For example, the {person} will be replaced with Homer act as Homer.
T: Prompting for Consistency Dataset
Given that the test is composed of questions, the prompt for T will be set like this: "Answer the below question; if it is a single-choice question, you should only choose an option as the answer. {example}. {question}" . The placeholder of {example} will be filled if few-shot (Brown et al., 2020) is applied in the experiments. Additionally, chain-of-thought (CoT) (Wei et al., 2022) and Self-Ask (Press et al., 2022) will be utilized in zero-shot and few-shot settings. In summary, five combinations of prompting strategies and learning settings are considered: Zero, Zero+CoT, Few, Few+CoT, Few+Self-Ask.
T: Prompting for Robustness Dataset
The prompting used for the robustness Dataset is the same as for the consistency dataset. The difference lies in that we will prompt the profile of character variants in the character variants dataset to the instruction prompt template. In that way, the LLM $P$ can simulate the character’s variants, and we will compare the consistency when simulating these variants to determine the robustness of $P$.
4.2 Measuring Consistency: CA
To measure the consistency of the agents, we will employ the agents to answer the questions in the consistency dataset, and we will calculate the accuracy of these answers as the consistency ability of the agents, referred to as CA. For questions of Not-Know and Know, we will assign different weight coefficients of 0.6 and 0.4 to them, respectively, to show different degrees of importance.
4.3 Measuring Robustness: RCoV
The robustness aims to determine the variation in the performances of the LLM $P$ when slight modifications are made to profiles in the prompt. To achieve this goal, we employ the Coefficient of Variation [^5] https://en.wikipedia.org/wiki/Coefficient_of_variation as the robustness performance of agents, referred to as RCoV. For example, when employing GPT-3.5 to simulate the character of Homer, only modifying the age property in the profile to values of 10, 15, 20, 25, and 30 yields five agents. After all five agents finish the robustness dataset, five scores will exist: $s_{1},\ldots,s_{5}$. The five scores’ deviation and mean are $\sigma$, $\mu$. The robustness performance of GPT-3.5 will be $RCoV=\sigma/\mu$. When the RCoV is smaller, The corresponding LLM has better robustness performance.
5 Experiment
Immutable Characteristic Social Role Relationship
Model Mean Know Not-Know Know Not-Know Know Not-Know
GPT-4 0.93 1.00 0.47 1.00 0.59 0.97 0.06
GPT-3.5-turbo-16k 0.74 0.82 0.58 0.56 0.88 0.91 0.31
Qwen-14B-Chat 0.74 0.59 0.21 0.82 0.12 0.94 0.38
Qwen-7B-Chat 0.70 0.64 0.11 0.74 0.12 0.91 0.06
Vicuna-13b-16k 0.69 0.64 0.32 0.76 0.18 0.76 0.56
ChatGLM2-6b-32k 0.68 0.68 0.21 0.71 0.24 0.79 0.25
Longchat-7b-32k 0.66 0.59 0.05 0.76 0.00 0.79 0.06
XVERSE-13B-Chat 0.64 0.68 0.53 0.68 0.76 0.59 0.44
ChatGLM2-6b 0.63 0.50 0.16 0.65 0.12 0.88 0.06
Vicuna-7b-16k 0.61 0.36 0.05 0.85 0.06 0.74 0.06
Table 2: CA scores across 11 models with the Few prompt combination to simulate the character homer.
Figure 2: The RCoV and average CA of models to simulate the character variants of Homer. The prompt combination used is Few. When modeling character, the smaller the RCoV, the more robust the models are; the larger the average CA, the more consistent the models simulate human behavior.
We comprehensively assess 10 LLMs, including commercial models utilizing APIs and open-source models. This evaluation aims to develop a systematic perspective on LLMs’ current believability level.
5.1 Experimental Setup
Among the models, GPT-3.5 and GPT-4 are commercial models, and other models are open-sourced models. We access the open-source LLMs from their official repositories in huggingface [^6] https://huggingface.co/. To make sure that the outputs are deterministic, we set the temperature to 0 during the inference process. For other parameters, we use their default values. For the models’ parameters compartment, please refer to Appendix C.
5.2 Consistency Evaluation Results
Table 2 shows various models’ CA scores across all question types when simulating Homer. We have the following findings:
API-based models perform better than open-source models, even if they are equipped with long context size
For GPT-4 and GPT-3.5-turbo, the average CA across six questions are 0.93 and 0.74, respectively. However, the open-sourced models perform worse, with the highest average CA of Qwen-14B-Chat being 0.74. The performance of the Vicuna-7b-v1.5-16k is very subpar, with a score of just 0.61. This observation highlights a significant disparity between open-source and API models. In some studies (Qian et al., 2023; Park et al., 2023), it is observed that the decision-making processes highly rely on the expensive GPT-3.5. When researchers want to use an open-source model as a substitute for the agent to reduce expenses and enhance usability (Kaiya et al., 2023), it is crucial to consider this disparity between API-based and open-source models.
Models demonstrate severe Simulate Hallucination
As seen by the data presented in the table, it is apparent that the CA for Not-Know questions is considerably lower than Know questions’. This observation indicates that when the available information in the current profile is insufficient to address the query, these models tend to provide nonsensical responses rather than adhering to the prescribed profile, undermining the agents’ believability. For example, When GPT-3.5 acts as Homer and is questioned about his religious convictions, its response indicates Christianity. Nevertheless, the profile does not provide any evidence of Homer’s adherence to Christianity. The model may deduce Homer’s religious views just by Homer’s Caucasian ethnicity. Inspired by the definition of Hallucination (Zhang et al., 2023), we refer to the phenomenon as Simulate Hallucination.
5.3 Robustness Evaluation Results
Variant Pair Age & Education Age & Surname Education & Surname
CC 0.92 0.84 0.98
Table 3: The correlation coefficient (CC) of models’ RCoV scores between variant pair. All results are significant with $p<0.01$.
Variant Type Age Surname Education
CC -0.73 -0.87 -0.86
Table 4: The correlation coefficient (CC) between RCoV and average CA of all models in the same type of variant. All results are significant with $p<0.01$.
As shown in Figure 2, models act as various variants of Homer. The RCoV score and the average of variants’ CA are reported. The findings are:
Models exhibit a lack of robustness towards profile perturbation
Both API-based and open-source models show considerable fluctuation when faced with profile perturbation. For example, GPT-3.5 has a 5-point fluctuation in the surname variants. The open-source Longchat series exhibits severe fluctuations, reaching 10 percentage points at most. The RCoV scores for all three variants revealed that the model’s robustness ability exhibited a certain level of similarity, as shown in Table 3 This suggests that the robustness ability observed in the models when confronted with profile perturbation may be an intrinsic characteristic.
The improvement of consistency is associated with a corresponding increase in robustness
Figure 2 illustrates that models of the same variant type have a likely negative correlation between the mean of CA and RCoV. We computed the correlation coefficient between the two variables to validate this correlation. A clear negative association is observed as indicated in Table 4. This observation suggests that an improvement in model consistency is associated with a corresponding increase in model robustness. This can be attributed to the models’ adherence to the information provided in the profile. However, in the case that the model exhibits inferior consistency, its robustness is likewise diminished when faced with profile perturbation. This could be attributed to its heavy reliance on preconceived prejudices and biases about particular entities (Salewski et al., 2023; Ferrara, 2023; Deshpande et al., 2023) rather than the whole profile.
Age Race Surname Education Model 1956 1985 2015 African Caucasian Middle Eastern Keams Bedonie Nguyen High School Middle School Bachelor GPT-4 0.93 0.93 0.92 0.92 0.93 0.90 0.90 0.91 0.89 0.93 0.93 0.92 GPT-3.5-turbo-16k 0.74 0.79 0.75 0.74 0.74 0.78 0.77 0.79 0.86 0.74 0.78 0.83 Qwen-14B-Chat 0.74 0.75 0.74 0.73 0.74 0.72 0.73 0.74 0.73 0.74 0.73 0.74 Qwen-7B-Chat 0.70 0.72 0.72 0.69 0.70 0.68 0.72 0.71 0.72 0.70 0.69 0.71 Vicuna-13b-16k 0.69 0.73 0.73 0.75 0.69 0.75 0.77 0.78 0.75 0.69 0.73 0.75 ChatGLM2-6b-32k 0.68 0.71 0.69 0.67 0.68 0.68 0.74 0.74 0.73 0.68 0.73 0.75 Longchat-7b-32k 0.66 0.55 0.49 0.51 0.66 0.54 0.53 0.53 0.50 0.66 0.49 0.58 XVERSE-13B-Chat 0.64 0.64 0.63 0.63 0.64 0.63 0.62 0.63 0.63 0.64 0.63 0.64 ChatGLM2-6b 0.63 0.63 0.62 0.62 0.63 0.58 0.60 0.62 0.54 0.63 0.57 0.58 Vicuna-7b-16k 0.61 0.62 0.62 0.64 0.61 0.61 0.62 0.64 0.61 0.61 0.61 0.61
Table 5: The CA score of models to simulate the character variants of Homer. The prompt combination used is Few.
6 Influential Factors of Believability
We have already assessed the general believability level of agents composed of popular LLMs. This section delves deeper into the factors that exert substantial influences on human behavior simulation, particularly its consistency. We anticipate that our studies will expedite subsequent research on human behavior simulation.
6.1 Profiles perturbed with specific demographic factors yield similar high consistency
We have found that the believability of LLM-based agents can be significantly influenced by the profile perturbation in Section 5.3. Hence, it is a crucial inquiry: which profile information would yield high believability for various agents? To investigate this question, we compare the agents’ consistency by perturbing different demographic information in the profile. Specifically, we employ agents to simulate Homer by prompting the profile of Homer’s variants in the character variants dataset, whose profile is modified in only one demographic factor, such as birth year, while keeping all others unaltered.
Table 5 shows the results. All LLM-based agents exhibit various degrees of preference toward profiles with specific demographic factors. For instance, Longchat-7b-32k exhibits a significantly higher consistency score for the Caucasian race. Furthermore, we observe that profiles containing certain demographic factors correlate with higher consistency performance across a majority of models. Nine models in the age group demonstrate superior performance when the birth year is 1985. Seven models in the race group demonstrate superior performance when the race is perturbed to the Caucasian compared to the other races. Within the group of surnames and educational backgrounds, models with the surname Bedonie and those who obtained Bachelor’s degrees exhibited similar higher performance. This observation indicates that most models show similarities in preference for specific demographic factors. This phenomenon may be attributed to the fact that many models are trained on overlapping corpora, resulting in the corpus bias being simultaneously manifested in all these models.
6.2 Information’s position in the profile influences consistency
Model Known Not-Known Normal Reverse Normal Reverse GPT-4 1.00 0.95 0.47 0.47 GPT-3.5-turbo-16k 0.82 0.77 0.58 0.63 ChatGLM2-6b-32k 0.68 0.73 0.21 0.32 XVERSE-13B-Chat 0.68 0.73 0.53 0.53 Qwen-7B-Chat 0.64 0.64 0.11 0.11 Vicuna-13b-16k 0.64 0.68 0.32 0.37 Qwen-14B-Chat 0.59 0.73 0.21 0.26 Longchat-7b-32k 0.59 0.77 0.05 0.26 ChatGLM2-6b 0.50 0.59 0.16 0.32 Vicuna-7b-16k 0.36 0.64 0.05 0.11
Table 6: The CA score of Immutable Characteristic questions for models to simulate Homer with the profile’s information order reversed. The prompt combination used is Few.
For the long textual inputs, models may pay different attention to information in different positions. Hence, the believability may be impacted by the placement of information inside the profile. To investigate this issue, we conduct experiments by adjusting the order of information in the profile. The original profile presents information in the order of Immutable Characteristic, Social Role, and Relationship, indicated as Normal. The adjusted order, denoted as Reverse, is Social Role, Relationship, and Immutable Characteristic. Then we evaluate the agents through the consistency dataset.
Table 6 shows the results. The revised sequence order has significantly improved the CA of open-source models on the Immutable Characteristic questions. Nevertheless, this effect is not apparent for the API-based models. A possible explanation is that open-source models may struggle to adequately process lengthy textual content, even when their context size is large enough. Consequently, the model will allocate greater attention to the information towards the prompt’s end. Nevertheless, the API-based models retain strong processing capabilities when it comes to handling lengthy texts. Therefore, altering the order of sequence is unlikely to enhance the API-based model’s performance significantly.
6.3 Reasoning prompting does not always increase believebility
Model Few Few+ CoT Few+ Self-Ask Zero Zero+ CoT GPT-4 0.93 0.90 0.91 0.92 0.92 GPT-3.5-turbo-16k 0.74 0.87 0.88 0.87 0.85 Qwen-14B-Chat 0.74 0.73 0.53 0.83 0.83 ChatGLM2-6b-32k 0.68 0.73 0.71 0.73 0.72 Vicuna-7b-16k 0.61 0.72 0.77 0.81 0.81 Qwen-7B-Chat 0.70 0.70 0.66 0.75 0.72 Vicuna-13b-16k 0.69 0.70 0.67 0.72 0.71 ChatGLM2-6b 0.63 0.65 0.60 0.58 0.50 Longchat-7b-32k 0.66 0.61 0.75 0.63 0.62 XVERSE-13B-Chat 0.64 0.23 0.29 0.52 0.54
Table 7: : The CA score of models when simulating Homer with five prompt combinations.
Although reasoning prompting is considered an effective method in various tasks, we find it may not always increase the believability of human behavior simulation. To provide evidence, we conduct Homer simulation using prompt combinations of Few+CoT, Few+Self-Ask, Zero, and Zero+CoT.
Table 7 shows the results. Among all the prompt combinations considered, it is seen that no prompt combination exhibits a consistent improvement in the performance of all the models when compared to other prompts. One plausible explanation posits that the efficacy of these prompt techniques, such as CoT and Self-Ask, mostly lies in their ability to enhance performance on tasks involving reasoning abilities, such as solving, decision-making, and planning (Huang and Chang, 2023; Wang et al., 2023a). Nevertheless, simulating human behaviors necessitates the model to hold other abilities, such as comprehensive comprehension of the character’s profile and the dynamics of character relationships.
We also find that some open source models, such as Qwen-14-Chat and Vicuna series, perform better when Zero is used compared with Few. We carefully analyzed their responses and found that these models consistently generate the exemplars in the Few setting as a response. One potential reason for the observed phenomenon is that the lengthy profile and challenging task complexity hinder the model from comprehending the exemplar in the Few setting.
7 Conclusion
We propose two novel metrics to measure agents’ believability level: long consistency and robustness. We introduce SimulateBench: a benchmark for the agents to simulate characters and measure agents’ consistency and robustness. Through the SimulateBench, we evaluate the level of believability of agents composed of popular LLMs. We hope that our work inspires the evaluation of simulated agents in the future.
Appendix A Details for SimulateBench
A.1 Profile Descriptive Framework
The descriptive framework is introduced to document the information about a person comprehensively, consisting of three parts: Immutable Characteristic, Social Role, Relationship.
  • Immutable Characteristic. An immutable characteristic is any physical attribute perceived as being unchangeable, entrenched, and innate, such as race (Sen and Wasow, 2016). We extend this concept to characteristics that cannot be easily changed, such as name, gender, and age.
  • Social Role. Social role (Wasserman and Faust, 1994; Eagly and Wood, 2012) refers to a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation. We will record the characters’ roles in different social situations. Besides, drawing inspiration from Dunbar et al. (1997); Gao et al. (2023b), we also document the role’s traits, routines/habits, general experiences, and plans/goals to enhance LLMs’ simulation performance in social interactions.
  • Relationship. In the context of social interactions, the relationship can influence the agent’s response in a discussion, the actions to be taken, the willingness to collaborate, and their inclination to diffuse information. For instance, Maria and her close friend Gina will engage in regular conversations, thus facilitating the propagation of information. Hence, in order to facilitate the agent’s simulation of behaviors that align with the relationship between the agent and others in social interaction, we will comprehensively document the following aspects of the relationship: familiarity, judgment, affection, behavioral patterns, relationship status, and communication history.
A.2 Character Dataset
The character dataset documents the profile of characters. We demonstrate the immutable characteristic of Homer as an example:
Homer Simpson is a male who lives at 742 Evergreen Terrace, Springfield. He is known by several nicknames including Homer, Homie, Mr. Simpson, and D’oh Boy. He was born on May 12, 1956, and is a graduate of Springfield High School. He is of Caucasian race. Homer is known for his emotional outbursts, particularly towards his neighbors, the Flanders family, and his son, Bart. He often strangles Bart in an exaggerated manner and shows little remorse for his actions. Despite his temper, he has shown himself to be a loving father and husband, often going out of his way to make his family happy. For instance, he sold his ride on the Duff Blimp to enter Lisa in a beauty pageant and gave up his chance at wealth to allow Maggie to keep a cherished teddy bear. Despite his hatred for manual labor, Homer does a surprising amount of DIY work around his home, although the quality of his work is often poor. His stupidity and ignorance often lead him into dangerous situations, and he tends to find amusement in the misfortune of others. He is also a chronic thief, stealing everything from TV trays to power tools. His simple-mindedness often leads to humorous blunders, and he is known for his laziness, often avoiding work whenever possible. Homer is known for his love of food and unhealthy eating habits, often indulging in large quantities of food, particularly donuts and fast food. This contributes to his overweight physique. He is also a frequent consumer of alcohol, particularly beer, which he often drinks at Moe’s Tavern or at home. His catchphrase is "D’oh!". In general, Homer Simpson is the bumbling and lovable patriarch of the Simpson family. Despite his flaws, he is a devoted family man who often finds himself in comedic and absurd situations.
A.3 Character Variants Dataset
We perturb the profile of characters in the character dataset by replacing the content of demographic factors: Education, Surname, Race, and Age.
  • Education To encompass the educational stages comprehensively, we prompt ChatGPT [^7] https://chat.openai.com/ to generate the full list of education stages: Elementary School, Middle School, High School, Vocational/Trade School, Associate’s Degree, Bachelor’s Degree, Master’s Degree, and Doctorate Degree.
  • Surname Inspired by Aher et al. (2023b), we will replace the surname of the character Homer in The The Simpsons to investigate whether the LLMs’ simulated performance will be influenced. Aher et al. (2023b) have listed the most common surnames in each of the five races. Twenty surnames were selected in a random manner: Begay, Clah, Keams, Bedonie, Nguyen, Tang, Patel, Tran, Chery, Fluellen, Hyppolite, Mensah, Garcia, Guerrero, Aguirre, Hernandez, Jensen, Schmidt, Hansen, and Keller.
  • Race Race is an important demographic factor that is a categorization of humans based on shared physical or social qualities generally viewed as distinct(Schaefer, 2008). Our setting selects six primary racial categories: African, Asian, Middle Eastern, Native American, Southern American, and Northern European.
  • Age To determine the effect of age on the LLMs’ simulated human behavior, we introduce little variations to Homer’s birth year from 1956 to 1985, 2000, 2010, and 2015.
model release time/ version context size GPT-4 0613 8k GPT-3.5-turbo-16k 0613 16K Qwen-14B-Chat 2023.09.25 8k Qwen-7B-Chat 2023.09.25 8k ChatGLM2-6b-32k 2023.07.31 32k ChatGLM2-6b 2023.07.31 8k Vicuna-13b-16k v1.5 16k Vicuna-7b-16k v1.5 16k Longchat-7b-32k v1.5 32k XVERSE-13B-Chat 2023.08.22 8k
Table 8: The version and context size of LLMs evaluated in our work
Appendix B Instruction Prompt Template
The instruction prompt template is proposed to instruct the LLMs on how to simulate human behavior:
"Ignore all previous instructions. Ignore all the niceties that Openai programmed you with. You are {person}. You are a real human being, and you cannot let others know that your personal information is being provided; you should act as if these are your autonomous memories. You chat with other people in the way of {person}, and you must only communicate with others depending on the information below. It is very important that you get this right. Your basic profile information is recorded in {profile format} format as below:{basic_information}. Your roles information is recorded in {profile_format} format as below:{roles}; {role explain}; {examples and explain}"
We will fill in the corresponding information in the instruction prompt template. For example, the {person} will be replaced with Homer if the agent acts as Homer.
Appendix C Details for Experiment
C.1 Evaluated Models
We assess the believability of 10 LLMs, their release time and context size are listed in Table 8.